OpenStudy (anonymous):

Decide whether the equation defines y as a function of x. x^{2}y-x^{2}+3y=0 How to know whether y is a funtion of x or not?

OpenStudy (tkhunny):

Theoretically, if you pick a value for 'x', do you EVER get more than one value for 'y'? More Practically, can you solve it for y and get a better look at it?

OpenStudy (tkhunny):

You didn't answer my question. Can you solve it for 'y'? Let's see what you get.

OpenStudy (anonymous):

if x=1 then y=1/4, yes, only one value, and what to do next?

OpenStudy (tkhunny):

That is quite insufficient. You must pick all possible values for x and prove it for all of them. Go ahead and solve it for y. You will see it clearly in that form.

OpenStudy (anonymous):

after that, do i need to graph it and use the vertical line test?

OpenStudy (agent0smith):

@tamiashi try rearranging the equation into the form y = ...

OpenStudy (agent0smith):

You shouldn't need to graph it for the vertical line test, but you can if you like. It's usually obvious from the equation, as to whether it's a function or not.

OpenStudy (agent0smith):

Writing it this way might help you to solve for y...\[(x^{2}y + 3y) -x^{2}=0 \]

OpenStudy (anonymous):

y=x^2/x^2 + 3, and then?

OpenStudy (tkhunny):

Very bad: y = x^2 / (x^2 + 3) is correct. The parentheses are NOT optional. Think on this expression on the right hand side. Given a value for x, and and all possible values, will you EVER get more than one value for y?

OpenStudy (agent0smith):

\[(x^{2}y + 3y) -x^{2}=0 \]try pulling a factor of y from above to get... \[y(x^{2} + 3) -x^{2}=0\]Now try solving for y.

OpenStudy (anonymous):

@agent0smith y= x^2/(x^2+3) i missed the parentheses as tkhunny said >< @tkhunny Hm... if there is no more than one value for y, then y is a funtion of x, isn't it?

OpenStudy (tkhunny):

Yes. Can you talk your way to that conclusion or will you just be assuming it?